PLEASE help me with this; I've been doing homework for hours and I don't know how to do this problem.
"After t seconds, the height (t) of a bottle rocket launched from the ground into the air is given by the function h(t) = -16tē + 80t. Find how long it takes the rocket to teach a height of 96 feet."
THANK YOU!

Answer by mananth(16949)

h(t) = -16tē + 80t.
h(t)=96 feet
find t
96 = -16t^2+80t
16t^2-80t+96=0
16t^2-32t-48t+96=0
16t(t-2)-48(t-2)=0
16t-48)(t-2)=0
16(t-3)(t-2)=0
t=3 & 2
It will take 2 seconds